OpenStudy (anonymous):

The distance from Kenya's house to her job is 18 miles. One morning she left her house at 7:15 a.m. She traveled at the rate of 45 miles per hour for 20 minutes, but to traffic she had to slow down to 20 miles per hour. At what time did Kenya arrive at work?

OpenStudy (aihberkhan):

Okay. To find how far she was at 45 miles per hour we would do \(20 \min = \frac{ 1 }{ 3 } hr\) \(\frac { 1 }{ 3 ]}* 45 = 15 mph\) So... \(15 - 15 = 3~miles~at~20~mph\) Now to find the time at \(20mph\) we do \(\frac { 3 }{ 20 } = .15hrs ~at ~20mph\) So the total time is \(.15 + .3333 = .4833 hrs\) or \(.4833 * 60 ~ 29 min\) 7:15 + :29 = 7:44 a.m. arrive at work

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @kittykatrave